OpenStudy (anonymous):

For the trinomial, determine the signs that will appear in the binomial factors. A. both + B. both - c. one +, one -

jimthompson5910 (jim_thompson5910):

The trinomial is missing

OpenStudy (anonymous):

x^2+740x+9810

OpenStudy (anonymous):

would the answer not be a?

jimthompson5910 (jim_thompson5910):

Do you know the quadratic formula?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

Use this formula to solve x^2+740x+9810 = 0 for x

jimthompson5910 (jim_thompson5910):

Once you get the solutions, you can use them to factor x^2+740x+9810

OpenStudy (anonymous):

ok so what do i plug in for b ?

OpenStudy (anonymous):

im kinda lost still

jimthompson5910 (jim_thompson5910):

a = 1, b = 740, c = 9810

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(740)\pm\sqrt{(740)^2-4(1)(9810)}}{2(1)}\]

OpenStudy (anonymous):

and the plus or minus sign what does that mean?

jimthompson5910 (jim_thompson5910):

Example: 10 plus/minus 5 is the same as 10+5 or 10-5

OpenStudy (anonymous):

im lost sorry

OpenStudy (anonymous):

i hate problems like these, and i don't have my caculator wwith me

jimthompson5910 (jim_thompson5910):

you can use google as a calculator

jimthompson5910 (jim_thompson5910):

or you can use wolfram alpha

jimthompson5910 (jim_thompson5910):

or you can use graphcalc http://www.graphcalc.com/

jimthompson5910 (jim_thompson5910):

There are a ton of online calculators or calculators you can download

OpenStudy (anonymous):

i believe it's going to be both +

jimthompson5910 (jim_thompson5910):

hmm are you sure that its x^2+740x+9810 that can't be factored

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that's the problem

jimthompson5910 (jim_thompson5910):

well the solutions are -370+sqrt(127090), -370-sqrt(127090) this means that it cannot be factored over the rationals

OpenStudy (anonymous):

so therefore it has to be both +'s?

jimthompson5910 (jim_thompson5910):

I'm not exactly sure what this problem wants then

OpenStudy (anonymous):

it just says determine the signss that will appear in the binomial factors

jimthompson5910 (jim_thompson5910):

i know

OpenStudy (anonymous):

so what would ur guess be?

jimthompson5910 (jim_thompson5910):

but it doesn't factor, unless they mean it factors to ( x+370-sqrt(127090) )( x+370+sqrt(127090) )

jimthompson5910 (jim_thompson5910):

but most books won't consider that factoring

OpenStudy (anonymous):

right so if it was u and the problem what would be ur guess as to what the signs would be

jimthompson5910 (jim_thompson5910):

hmm probably + and -, but that's just a wild guess

OpenStudy (anonymous):

it was both +'s that's what i thought 2

OpenStudy (anonymous):

im going to put another one up

jimthompson5910 (jim_thompson5910):

hmm then idk what they were after